OpenStudy (noah690):

Find f(–2) for f(x) = 2x - 5, is the answer -9. Thanks.

OpenStudy (trojanpoem):

f(-2) = 2 * -2 -5 = -4 - 5 = -9 ( +1 right)

OpenStudy (noah690):

I thought so, but it seemed too easy. Thanks.

OpenStudy (trojanpoem):

yw